# Vision-Based CNN Prediction of Sunspot Numbers from SDO/HMI Images

### v1

@@ 3.1 Model Interpretability
Figure 7 displays Grad-CAM overlays for representative days across different solar activity levels. The heatmaps highlight that the network focuses on localized dark sunspot regions as well as surrounding penumbral structures, consistent with human identification criteria. Importantly, even during periods of relatively low activity, the model’s attention remains concentrated on subtle intensity depressions corresponding to small sunspots or pores, suggesting that the CNN has learned meaningful solar morphology rather than spurious image artifacts.
Figure 7: Grad-CAM overlays for representative continuum solar images. Warmer colors indicate regions with higher attribution to the predicted sunspot number. The CNN predominantly attends to sunspot cores and penumbrae.
Complementary results from Integrated Gradients and SmoothGrad-IG are shown in Figure 8. Unlike Grad-CAM, which highlights localized regions, IG methods distribute importance scores more diffusely, often emphasizing granular textures across the photospheric background. This behavior suggests that the model leverages both the presence of discrete sunspots and the broader distribution of intensity patterns to form its regression output. The smoother attribution maps of SmoothGrad-IG further confirm the robustness of these relevance patterns.
Figure 8: Integrated Gradients (IG) and SmoothGrad-IG attribution overlays for representative samples. Columns show the original solar image, IG relevance map, and SmoothGrad-IG map. Results confirm that the CNN uses both sunspot regions and diffuse photospheric structures for prediction.
These interpretability analyses strengthen confidence in the physical plausibility of the model. Rather than relying on spurious correlations, the CNN consistently identifies sunspot-bearing regions as key drivers of prediction, aligning its internal representations with the established heliophysical understanding of sunspot number indices.
@@ 4 Discussion and Conclusions
This work demonstrates that convolutional neural networks trained directly on solar continuum images can provide highly accurate estimates of daily sunspot numbers. By avoiding explicit feature engineering or manual counting, the proposed approach leverages the full spatial information of solar imagery and achieves competitive performance with state-of-the-art methods in the literature. On the independent test set, our model reached $R^{2}=0.986$ and RMSE = 6.25, values that not only surpass traditional statistical approaches but also approach the accuracy of hybrid deep learning architectures reported in recent years.
Table 2 summarizes a selection of previous studies on sunspot number prediction using statistical, machine learning, and deep learning models. Results highlight three key trends: (i) classical methods such as ARIMA or ETS, while historically important, underperform in terms of error magnitude and explained variance; (ii) hybrid and ensemble models (e.g., SVM-LSTM or ARIMA-ETS-SVM) achieve higher accuracy but require complex pipelines and handcrafted features; (iii) deep learning methods, particularly convolutional and recurrent neural networks, consistently yield strong results, often exceeding $R^{2}>0.95$. This comparative perspective highlights that while hybrid models often achieve very high $R^{2}$, their complexity and dependence on handcrafted features contrast with the simplicity and transparency of our direct vision-based CNN.
Our CNN achieves performance comparable to or better than most hybrid pipelines, while maintaining a direct vision-based formulation that bypasses manual sunspot identification. Compared to hybrid SVM-LSTM models (e.g., Abdulkadir et al. 2023, $R^{2}=0.99$, RMSE = 7.77), our approach shows lower RMSE with slightly lower $R^{2}$, indicating a balanced trade-off between precision and generalization. Against Informer-based methods or HAVOK-ML (Yang et al. 2022), our model demonstrates a simpler but competitive alternative. Importantly, the ability to work directly on raw solar disk imagery makes this method attractive for operational monitoring, where interpretability via saliency methods (Grad-CAM, IG) can also guide scientific understanding.
@@
2025 CNN (vision-based, direct images) SDO/HMI + SILSO v2.0 0.986 6.25 This work
Future work should focus on three directions. First, combining raw imagery with historical time-series predictors may yield synergistic multimodal models that capture both morphological and temporal dynamics. Second, systematic benchmarking across entire solar cycles would provide robust comparisons against operational forecasts (e.g., NOAA, SIDC). Finally, integrating interpretability frameworks such as Grad-CAM and Integrated Gradients into operational monitoring may facilitate both scientific discovery and trust in automated forecasts.
This study demonstrates that vision-based regression is not only feasible but also highly effective for heliophysical index estimation, providing a foundation for the next generation of automated space weather monitoring systems.
Future developments should aim at multimodal architectures that combine imagery with time-series predictors, systematic benchmarks across cycles, and real-time operational deployment. By bridging physical interpretability with predictive accuracy, this study opens pathways toward next-generation automated space weather monitoring systems.

### v2

@@ 3.2 Model Interpretability
Figure 8 shows Grad-CAM overlays for representative dates spanning low to high solar activity. The highlighted regions concentrate on localized intensity depressions associated with sunspot umbrae and surrounding penumbrae, in agreement with the visual definition of sunspot-bearing structures in continuum images. Across all examples, the strongest attributions remain confined to compact active-region patches rather than the quiet photosphere or the image margins, supporting that the CNN relies primarily on physically meaningful morphology.
Figure 8: Grad-CAM overlays for representative HMI continuum quick-look images. Warmer colors indicate higher attribution to the predicted sunspot number. The examples correspond to (left-to-right, top-to-bottom): 2011–09–19, 2014–02–07, 2014–06–10, 2017–03–31, 2023–02–10, 2023–07–29, 2023–08–16, and 2023–09–22. In all cases, the strongest attribution concentrates on sunspot umbra/penumbra regions.
Complementary attribution maps from Integrated Gradients (IG) and SmoothGrad-IG are shown in Figure 9. Compared with Grad-CAM (which is tied to convolutional feature maps and thus tends to localize salient regions), IG-based methods operate at the pixel level and often yield more spatially distributed responses on textured solar backgrounds. In our case, SmoothGrad-IG reduces high-frequency noise and consistently reinforces the relevance of sunspot-bearing areas, while also indicating a weaker sensitivity to broader photometric context (e.g., global contrast variations) that can modulate regression outputs. Accordingly, we interpret IG/SmoothGrad-IG as a complementary, qualitative check rather than a precise segmentation of sunspot pixels.
Figure 9: Integrated Gradients (IG) and SmoothGrad-IG attribution overlays for representative samples. Columns show the original image, IG overlay, and SmoothGrad-IG overlay. IG-based methods provide complementary, pixel-level relevance patterns; SmoothGrad-IG reduces high-frequency noise and reinforces the contribution of sunspot-bearing regions while also reflecting weaker sensitivity to global photometric context.
We emphasize that these explainability tools serve different purposes: Grad-CAM provides coarse but spatially localized evidence tied to deep convolutional representations, whereas IG variants can appear more diffuse on textured inputs because they attribute relevance at the pixel level. Using both perspectives helps validate that the dominant signal remains aligned with sunspot structures.
Overall, the explainability results support the physical plausibility of the learned mapping: the dominant attribution consistently aligns with sunspot structures in continuum intensity images, which are the direct visual drivers of the SILSO sunspot number.
@@ 4 Discussion and Conclusions
This work demonstrates that convolutional neural networks trained directly on solar continuum images can provide accurate estimates of daily sunspot numbers. By avoiding explicit feature engineering or manual counting, the proposed approach leverages the full spatial information of solar imagery and achieves competitive performance in an image-to-scalar regression setting. On the independent test split, our model reached $R^{2}=0.964$ with RMSE =9.75=9.75 and MAE =6.74=6.74, indicating strong agreement with the SILSO daily index across a wide activity range.
Table 2 summarizes representative studies on sunspot-number modeling using statistical, machine learning, and deep learning approaches. The comparison should be interpreted with caution because many works address different targets (monthly means, smoothed indices, or multi-step estimation) and different input modalities (time series rather than images). Still, three broad trends emerge: (i) classical statistical forecasters (e.g., ARIMA/ETS variants) provide useful baselines but typically lag behind modern ML/DL methods in accuracy for long-horizon or multi-step settings; (ii) hybrid and ensemble pipelines can achieve high skill but often rely on decomposition steps and engineered representations; and (iii) deep learning models (convolutional and recurrent/attention-based) consistently report strong performance, especially on aggregated or smoothed sunspot indices. Within this landscape, our contribution is complementary: a direct, vision-based mapping from full-disk imagery to the daily sunspot number.
Our CNN provides a competitive baseline for direct image-based estimation while keeping the pipeline simple and fully end-to-end. Relative to image-based feature-engineering approaches (e.g., topological descriptors followed by classical regressors), the present model removes the need to define handcrafted image features and enables interpretability analyses directly on the learned representations. At the same time, we emphasize that many recent high-$R^{2}$ reports in the literature correspond to monthly or smoothed time-series forecasting tasks, which are not directly comparable to daily image-to-scalar estimation.
@@
2025 CNN (vision-based, direct images) SDO/HMI + SILSO v2.0 (daily) 0.964 9.75 This work
Future work should focus on three directions. First, a synergistic multimodal model would combine an image encoder (e.g., a CNN/ViT operating on HMI continuum images) with complementary temporal predictors (e.g., recent sunspot-number history and other scalar proxies such as F10.7), using feature fusion (early or late fusion) to capture both instantaneous morphology and short-term temporal context. Second, systematic benchmarking across cycles should be pursued via cross-cycle holdouts (e.g., training on Solar Cycle 24 and testing on Cycle 25, and vice versa) and rolling-origin evaluations, to quantify robustness under domain shifts and changing activity regimes. Third, interpretability tools (Grad-CAM, IG) can be expanded into routine quality-control diagnostics, supporting trust and scientific plausibility checks in automated monitoring pipelines.
In summary, our results show that direct, vision-based regression from full-disk continuum images to the daily SILSO sunspot number is feasible and accurate, providing a practical baseline for integrating modern computer vision into automated sunspot-number estimation.